OpenStudy (anonymous):

if cos(pi*n) = (1)^n --- what is cos ([pi*n]/4)?

OpenStudy (anonymous):

its me again and i want to explain some more, you see example is sin ( pi*n/2) = 1/2* i * [ (-i)^n - (i)^n ] try it because sin (pi*n / 2 ) goes {0,1,-1,0} cos ( pi * n / 2 ) = {almost the same } 1/2 *i* [ (-i)^(n+1) - (i)^(n+1) ] now does anybody understand to help me get the cos ([pi*n]/4)??

OpenStudy (anonymous):

more information about cos ([pi*n]/4) is that is for n equals en interger from 0 to 8 {1, sqrt(2)/2 , 0 , -sqrt(2)/2 , -1 , -sqrt(2)/2 , 0 , sqrt(2) , 0 , 1}